Find the slope of the line described by –3x + 5y = –30.
Answer by Maths68(1474)

Find the slope of the line described by –3x + 5y = –30.
Equation of the line slope-intercept form
y=mx+b
m=slope
b=y-intercept
Rearrage the given equation accroding to the slope-intercept form
-3x+5y=-30
5y=3x-30
Divide by 5 both sides of the equation
5y/5=(3x-30)/5
y=3x/5-30/5
y=3/5(x)-6
Compare the above equation with Equation of the line slope-intercept form
Slope=m=3/5
y-intercept = b = -6